A spherical storage tank has a diameter of 48 ft and is full of water. How many delivery trucks with cylindrical tanks measuring 8 ft in diameter and 10 ft long can be completely filled from the storage tank?

A)

230 trucks

B)

115 trucks

C)

86 trucks

D)

1152 trucks

Answer by checkley77(12844)

V=4/3PI*R^3
V=4/3*3.14*24^3
V=4.187*13,824
V=57,881.09 TANK VOLUME.
V=PI*R^2*H
V=3.14*4^2*10
V=3.14*16^10
V=3.14*160
V=502.4 VOLUME OF THE TANKER TRUCK.
57,881.09/502.4=115.21 TANKER TRUCK LOADS ARE REQUIRED TO EMPTY THE TANK.
ANSWER B)